# Ditch the Denoiser: Emergence of Noise Robustness in Self-Supervised Learning from Data Curriculum

### v1

@@ 1 Introduction
This is especially common in medical imaging [27, 42, 26], astrophysics [29], geological rock imaging [6], remote sensing [14] and finance [28]. For example, the earth radar images taken by ESA Sentinel-1 [14] comes with speckle noise, and the clean data for supervised denoising is difficult to acquire since it requires better sensor and atmospheric conditions. As another example, Sloan Digital Sky Survey (SDSS) imaging in the COSMOS region contains significant sky noise [29] which can obscure faint celestial objects. Despite advancements in SSL, methods like DINOv2, as shown in Figure 5, still yield significant reduction in performance when trained directly on corrupted datasets.
To address the challenges outlined above, a trivial baseline is a denoiser-preprocessed pipeline, which trains a self-supervised denoiser on the noisy data first, then trains the SSL models on the denoised images. In downstream tasks, the SSL denoiser always serves as a preprocessor because the SSL model is only trained on denoised data thus cannot handle noisy inputs. While effective, this pipeline adds substantial inference latency, fine-tuning overhead, and deployment complexity (Figure 1, left). Instead, we explore training strategies that enable an off-the-shelf SSL backbone (e.g., DINOv2) to ingest noisy inputs directly, thereby ditching the denoiser during downstream use (Figure 1, right). To the best of our knowledge, this problem has little-to-no prior work in the existing literature. We take a first step by corrupting ImageNet [12] with synthetic noise that mirrors real-world degradations, providing a controlled yet challenging testbed for systematic evaluation and benchmarking.
@@
3. We introduce a new training paradigm for enabling models like DINOv2 to learn robust representations directly from noisy pretraining data. By combining noise curriculum learning with a novel denoised regularization loss, our method achieves fully denoiser-free downstream fine-tuning and inference while maintaining strong performance across tasks. As shown in Figure 2, our methods, DINOv2 w/ NC and DINOv2 w/ NCT improve significantly over the DINOv2 baseline, and closely matches the denoiser-preprocessed pipeline N2N + DINOv2.
@@ 3.1 Preliminary: Curriculum Learning Background and Denoiser-Preprocessed Baseline
[…]
@@ 3.2 DINOv2 w/ NC (Noise Curriculum): Curriculum Learning for Self-Supervised Representation Learning
Ideally, instead of relying on a separate denoiser, the representation model should develop robustness to noise on its own, enabling an end-to-end training setup in downstream tasks. Figure 3 outlines our curriculum learning approach to achieve this. Following the definition in Section 3.1, we design a curriculum $\langle Q_{1},Q_{2}\rangle$ that satisfies the key requirement $H(Q_{1})<H(Q_{2})$. Here, $Q_{1}$ corresponds to the distribution over denoised images $X_{\text{denoised}}$, which lies in a more structured, lower-entropy subspace. $Q_{2}$ corresponds to the distribution over the original noisy images $X$, which has higher entropy due to the injected stochastic corruption that expands the support of the data distribution.
Such a curriculum is further motivated by the following toy example. We consider a simple MLP with 1 hidden layer trained on MNIST [22] using ReLU activation, and optimize the following self-supervised joint embedding loss Equation 1 that applies a soft identity constraint on empirical covariance matrix, which is a canonical formulation that aligns embeddings of different augmented views while preventing representational collapse:
@@
where $\hat{\Sigma}_{\theta}=\frac{1}{n}\sum_{i=1}^{n}(z_{i}-\bar{z})(z_{i}-\bar{z})^{\top},z_{i}=m_{\theta}(\tau(x_{i}))$, $m$ is the MLP and $\tau$ is the augmentation. Under Gaussian noise ($\sigma=0.4\cdot 255$), we observed that: Training and linear probing on the clean train set (50 epochs each) yields $91.21\pm 0.36\%$ (1-sigma) accuracy on the clean test set. Training and linear probing on the noisy train set (50 epochs each) yields $64.55\pm 3.47\%$ on the noisy test set. Surprisingly, curriculum learning (training 30 epochs on clean, then 20 on noisy), followed by 50 epochs of linear probing on noisy, recovers performance to $83.05\pm 0.45\%$ on the noisy test set. This shows that the model can adapt to the noisy data without forgetting what it learned from the clean phase. The significant recovery in performance (see Appendix Figure 6), even with a simple MLP and vector data, suggests that curriculum learning encourages the model to internalize noise-robust features. We now build on this intuition to scale our method to complex vision models like DINOv2:
@@
Denoiser-free Downstream Tasks and Inference: For downstream tasks involving noisy images, such as classification, the representations learned by $g_{\theta}$ can be directly utilized with a prediction head $h_{\theta}$ for task-specific learning. Further fine-tuning can be performed directly on the noisy data $z$ without requiring a denoising module, since the training strategy internalizes noise-robust features into the learned representations. This enables the entire model, both backbone $g_{\theta}$ and head $h_{\theta}$, to be optimized end-to-end on the downstream objective and deployed without any preprocessing at inference time as shown in Equation 2:
[…]
@@
Before restarting the noisy training, we first extract the weights of the teacher backbone $T_{\mathrm{dn}}$ that was trained on denoised data and freeze it. As shown in Figure 4, the architecture of the model now contains three components: teacher $T$, student $S$ and frozen denoised teacher $T_{\mathrm{dn}}$. When restarting the training, given a noisy image $x$ and its denoised version $x_{\mathrm{dn}}=f_{\theta}(x)$, we apply exactly identical augmentations (i.e., same crops, blur, flip…) to $x$ and $x_{\mathrm{dn}}$, so their embeddings are aligned. In addition to the original DINOv2 loss, we encourage the output scores of the student to be close to that of the denoised teacher. This prevents the training from deviating significantly from its cleaner initialization. So the denoised-regularized loss is given as:
@@
It is important to emphasize that this regularization technique is specifically designed to complement our curriculum training pipeline. For the regularization to be effective, the frozen teacher $T_{\mathrm{dn}}$ and the trainable teacher $T$ must have alignment in their output embeddings at the beginning of the restart training. This is accomplished by assigning them identical weights. Using a different frozen teacher, even if trained on the same denoised dataset, would not provide meaningful regularization due to misalignments, as shown in Appendix Figure 9. Like NC, NCT is also denoiser-free at downstream fine-tuning and inference.
@@ 4.1 Practical and Challenging Experimental Settings
Noise Addition. We follow ImageNet-C [18] to introduce three types of noise: Gaussian, Shot (Poisson), and Speckle noise, as these are among the most commonly encountered noise types in natural imaging. For the ImageNet-100 experiments, we use Gaussian noise ($\sigma=50,100,255$; SNR: 8.84 dB, 4.09 dB, 0.43 dB), Shot noise ($\lambda=10,3,1$; SNR: 8.55 dB, 4.16 dB, 0.58 dB), and Speckle noise ($\sigma=102,178.5,255$; SNR: 9.17 dB, 5.59 dB, 4.00 dB). For the ImageNet-1k experiments, we tested Gaussian noise at $\sigma=100$ and $\sigma=255$, with corresponding SNRs of 4.36 dB and 0.72 dB respectively. These SNRs are common values reported in real noisy medical imaging [40]. The example visualizations for each noise level are provided in Section C.2. The noises are added to raw images before preprocessing (e.g., resizing) to better reflect practical scenarios such as device or environmental noise. Consequently, while the initial noise level is fixed, the preprocessing steps introduce variations in actual noise levels that challenge the model’s ability to adapt effectively.
Implementation Details. In ImageNet-100 experiments, we use ViT-S/16 as the base architecture for DINOv2, and unless otherwise mentioned, train it for 200 epochs with a batch size of 40 using the AdamW optimizer. The base learning rate is $7.9\times 10^{-4}$ (scaled by the square root of batch size). The denoiser is trained for 100 epochs with a batch size of 4 using Adam. The linear probe uses a batch size of 128 and is trained for 12.5k steps. In ImageNet-1k, we adopt ViT-B/16 as the backbone and train DINOv2 for 100 epochs with a batch size of 512 and drop path rate of 0.1. The learning rate is $2.8\times 10^{-3}$. The denoiser is trained on a 100k subset with a batch size of 8 for 50 epochs (for Gaussian $\sigma=255$) and 90 epochs (for Gaussian $\sigma=100$). The linear probe uses batch size 1024 for 25k steps. Each training is done on a single RTX 4090 (~8h) or 4$\times$×L40S GPUs (~65h). The strength parameters are determined via light parameter sweeps with the details provided in Section A.2.
@@ 4.2 Linear Probing Reveals Superior Noise Adaptation
We begin by evaluating how different methods adapt to various levels and types of noise using a 200-epoch training schedule on ImageNet-100. We train a linear classifier over a frozen DINOv2 backbone on the training set, and report classification performance over the validation set. We compare the classification performance of the DINOv2 backbone over four training strategies. (1) N2N + DINOv2: train the backbone on denoised images for 200 epochs and evaluate the model on denoised images (2) DINOv2 w/ NC: train the backbone on denoised images for 140 epochs, then restart training on noisy images for 60 epochs without regularization. Finally, evaluate the model on noisy images. The rationale for such partitions of epochs is provided in Appendix B.3. (3) DINOv2 w/ NCT: same as the DINOv2 w/ NC but with the regularization loss applied in the 60 epochs of noisy training. We only evaluate the NCT method on the four most noisy scenarios (lowest SNRs), i.e., Gaussian noise ($\sigma$=255, $\sigma$=100), Shot noise ($\lambda$=1), and Speckle noise ($\sigma$=255). (4) DINOv2: train the backbone on noisy images for 200 epochs, and evaluate the model on noisy images. Additionally, we train and evaluate the model on original clean images, referred to as Clean. Figure 5 illustrates the classification accuracy curves.
As shown in Figure 5, our proposed method (green) significantly outperforms the backbone trained solely on noisy images (red) across all noise types and noise levels. At moderate noise levels, i.e., the first two columns, the accuracy of our method closely matches and even exceeds the backbone that is trained and evaluated on explicitly denoised data (shown in orange lines). Only at extreme noise levels, i.e., the last column, the gap between DINOv2 w/ NC and N2N + DINOv2 widens, primarily due to the significant loss of the original image signal. For instance, as shown in Section C.2, with Gaussian noise at $\sigma=255$, many images become virtually unrecognizable to the naked eye. Hence, regularization is introduced to guide the model in these extremely noisy scenarios. As illustrated in Figure 5, regularization DINOv2 w/ NCT improves the model’s performance at all four scenarios. The improvement amplifies as the gap between N2N + DINOv2 and DINOv2 w/ NC widens, with a substantial improvement of 6.7% over NC achieved at Gaussian noise with $\sigma=255$. However, when N2N + DINOv2 is on par with DINOv2 w/ NC, the regularization provides diminishing benefit, as demonstrated with Gaussian noise at $\sigma=100$. This is expected since the guidance becomes redundant when its representation quality is comparable to or lower than that of the original model.
@@
(a) Evolution of linear probing accuracies on ImageNet-100 under extended training durations. Relative ranking among methods remain consistent across epochs. See visualization in Figure 2.
Method Clean Gaussian-100 SNR: 4.36 dB Gaussian-255 SNR: 0.72 dB
@@ 4.3 Scaling Improves Performance Across Datasets
Longer Training Duration. We further improve the performance of our methods by extending the training on ImageNet-100 to longer epochs. For 500-epoch configuration, we restart the training from the 200-epoch denoised checkpoint (i.e., N2N + DINOv2) as described in Section 4.2; for the 1000-epoch configuration, we restart the training from the 500-epoch denoised checkpoint. As shown in Table 1(a) and Figure 2, accuracies increase as epoch increases and the relative ranking among methods remains consistent across training durations: N2N + DINOv2 achieves the highest accuracy, followed closely by DINOv2 w/ NC or NCT, while the baseline DINOv2 remains the lowest. This demonstrates the scalability of the proposed methods. Interestingly, we observe that DINOv2 w/ NCT consistently converges to or outperforms its anchor’s accuracy, as indicated by $\dagger$ and $\ddagger$ in Table 1(a), demonstrating that the regularization objective successfully guides the model toward robust representations. Notably, the performance gap between DINOv2 and the other methods narrows as training progresses, suggesting that DINOv2 is sensitive to noise and benefits significantly from longer training. In contrast, our noise curriculum and regularization approach greatly accelerates convergence, achieving similar or better performance in roughly half the training time.
@@
Larger Dataset. In our controlled ImageNet-1k experiment, we constrain the DINOv2 backbone to undergo a total of 100 training epochs to enable fair comparison across methods. DINOv2 w/ NC: train the backbone on denoised images for 30 epochs followed by 70 epochs of training on noisy images. DINOv2 w/ NCT: While we find that restarting at 30 epochs yields the best performance for DINOv2 w/ NC, the denoised teacher at this early stage lacks sufficiently strong representations to serve as an effective regularizer. To address this, we continue training the denoised teacher to 100 epochs, then freeze it for regularizing the DINOv2 backbone for 70 epochs of noisy training. Since the frozen teacher and the DINOv2 backbone are derived from the same training run and pass through the same 30-epoch state, alignment is preserved. This allows us to meet the 100-epoch constraint while still leveraging a stronger teacher. Table 1(b) shows that DINOv2 w/ NCT improves over the DINOv2 baseline with substantial absolute gains of $1.4\%$ and $4.8\%$ at noise levels $\sigma=100$ and $\sigma=255$, respectively. DINOv2 w/ NC also achieves a strong improvement of $2.5\%$ under the more challenging $\sigma=255$ setting. The relatively smaller gain of DINOv2 w/ NC at $\sigma=100$ is due to the fact that, in large-scale settings with moderate noise, longer training alone significantly improves DINOv2’s robustness. Since DINOv2 w/ NC allocates a portion of the training budget to denoised pretraining, it shortens the noisy training phase, slightly limiting its benefit. Overall, while scaling training on large datasets improves DINOv2’s noise adaptation under moderate noise, denoised pretraining and regularization remain essential under high noise to maintain strong performance.

### v2

@@ 1 Introduction
This is especially common in medical imaging [28, 43, 27], astrophysics [30], geological rock imaging [6], remote sensing [17] and finance [29]. For example, the earth radar images taken by ESA Sentinel-1 [14] comes with speckle noise, and the clean data for supervised denoising is difficult to acquire since it requires better sensor and atmospheric conditions. As another example, Sloan Digital Sky Survey (SDSS) imaging in the COSMOS region contains significant sky noise [30] which can obscure faint celestial objects. Despite advancements in SSL, methods like DINOv2, as shown in Figure˜5, still yield significant reduction in performance when trained directly on corrupted datasets.
To address the challenges outlined above, a trivial baseline is a denoiser-preprocessed pipeline, which trains a self-supervised denoiser on the noisy data first, then trains the SSL models on the denoised images. In downstream tasks, the SSL denoiser always serves as a preprocessor because the SSL model is only trained on denoised data thus cannot handle noisy inputs. While effective, this pipeline adds substantial inference latency, fine‑tuning overhead, and deployment complexity (Figure˜1, left). Instead, we explore training strategies that enable an off‑the‑shelf SSL backbone (e.g., DINOv2) to ingest noisy inputs directly, thereby ditching the denoiser during downstream use (Figure˜1, right). To the best of our knowledge, this problem has little-to-no prior work in the existing literature. We take a first step by corrupting ImageNet [12] with synthetic noise that mirrors real‑world degradations, providing a controlled yet challenging testbed for systematic evaluation and benchmarking.
@@
3. We introduce a new training paradigm for enabling models like DINOv2 to learn robust representations directly from noisy pretraining data. By combining noise curriculum learning with a novel denoised regularization loss, our method achieves fully denoiser-free downstream fine-tuning and inference while maintaining strong performance across tasks. As shown in Figure˜2, our methods, DINOv2 w/ NC and DINOv2 w/ NCT improve significantly over the DINOv2 baseline, and closely matches the denoiser-preprocessed pipeline N2N + DINOv2.
@@ 3.1 Preliminary: Curriculum Learning Background and Denoiser-Preprocessed Baseline
Curriculum learning refers to a training scheme that trains a machine learning model from easier to harder data that imitates the learning order in human curricula. This often leads to faster convergence and better performance. The seminal work [5] formally defines a curriculum as a sequence of training distributions $\langle Q_{1},...,Q_{t},...,Q_{T}\rangle$. Each distribution $Q_{t}$ is a reweighting of the target training distribution $P(z)$: $Q_{t}(z)\propto W_{t}(z)P(z)\ \ \forall z:\int Q_{t}(z)dz=1$ such that following conditions are satisfied: (1) the entropy of distributions increase over time, i.e., $H(Q_{t})<H(Q_{t+1})$. This implies that the model sees more diverse or complex examples as training progresses. A rising entropy indicates reduced certainty and increased difficulty. (2) Unnormalized weighting for any example increases over time, i.e., $W_{t}(z)\leq W_{t+1}(z)\ \forall z$. (3) Final distribution equals the target, i.e., $Q_{T}(z)=P(z)$. In practice, the first condition is most widely preserved, while the latter two are often relaxed to allow more flexible curricula [38]. In this work, we adopt this perspective and design our training schedule to follow an entropy-based curriculum from easy to hard examples.
@@ 3.2 DINOv2 w/ NC (Noise Curriculum): Curriculum Learning for Self-Supervised Representation Learning
Ideally, instead of relying on a separate denoiser, the representation model should develop robustness to noise on its own, enabling an end-to-end training setup in downstream tasks. Figure˜3 outlines our curriculum learning approach to achieve this. Following the definition in Section˜3.1, we design a curriculum $\langle Q_{1},Q_{2}\rangle$ that satisfies the key requirement $H(Q_{1})<H(Q_{2})$. Here, $Q_{1}$ corresponds to the distribution over denoised images $X_{\text{denoised}}$, which lies in a more structured, lower-entropy subspace. $Q_{2}$ corresponds to the distribution over the original noisy images $X$, which has higher entropy due to the injected stochastic corruption that expands the support of the data distribution.
Such a curriculum is further motivated by the following toy example. We consider a simple MLP with 1 hidden layer trained on MNIST [23] using ReLU activation, and optimize the following self-supervised joint embedding loss Equation˜1 that applies a soft identity constraint on empirical covariance matrix, which is a canonical formulation that aligns embeddings of different augmented views while preventing representational collapse:
@@
where $\hat{\Sigma}_{\theta}=\frac{1}{n}\sum_{i=1}^{n}(z_{i}-\bar{z})(z_{i}-\bar{z})^{\top},z_{i}=m_{\theta}(\tau(x_{i}))$, $m$ is the MLP and $\tau$ is the augmentation. Under Gaussian noise ($\sigma=0.4\cdot 255$), we observed that: Training and linear probing on the clean train set (50 epochs each) yields $91.21\pm 0.36\%$ (1-sigma) accuracy on the clean test set. Training and linear probing on the noisy train set (50 epochs each) yields $64.55\pm 3.47\%$ on the noisy test set. Surprisingly, curriculum learning (training 30 epochs on clean, then 20 on noisy), followed by 50 epochs of linear probing on noisy, recovers performance to $83.05\pm 0.45\%$ on the noisy test set. This shows that the model can adapt to the noisy data without forgetting what it learned from the clean phase. The significant recovery in performance (see Appendix Figure˜6), even with a simple MLP and vector data, suggests that curriculum learning encourages the model to internalize noise-robust features. We now build on this intuition to scale our method to complex vision models like DINOv2:
@@
Denoiser-free Downstream Tasks and Inference: For downstream tasks involving noisy images, such as classification, the representations learned by $g_{\theta}$ can be directly utilized with a prediction head $h_{\theta}$ for task-specific learning. Further fine-tuning can be performed directly on the noisy data $z$ without requiring a denoising module, since the training strategy internalizes noise-robust features into the learned representations. This enables the entire model, both backbone $g_{\theta}$ and head $h_{\theta}$, to be optimized end-to-end on the downstream objective and deployed without any preprocessing at inference time as shown in Equation˜2:
$\text{Finetuning:}\ \ \mathcal{L}=\mathbb{E}_{(z,y)}\left[L(h_{\theta}(g_{\theta}(z)),y)\right]\quad\text{Inference:}\ \ \hat{y}=h_{\theta}(g_{\theta}(z))$ (2)
@@
Before restarting the noisy training, we first extract the weights of the teacher backbone $T_{\mathrm{dn}}$ that was trained on denoised data and freeze it. As shown in Figure˜4, the architecture of the model now contains three components: teacher $T$, student $S$ and frozen denoised teacher $T_{\mathrm{dn}}$. When restarting the training, given a noisy image $x$ and its denoised version $x_{\mathrm{dn}}=f_{\theta}(x)$, we apply exactly identical augmentations (i.e., same crops, blur, flip…) to $x$ and $x_{\mathrm{dn}}$, so their embeddings are aligned. In addition to the original DINOv2 loss, we encourage the output scores of the student to be close to that of the denoised teacher. This prevents the training from deviating significantly from its cleaner initialization. So the denoised-regularized loss is given as:
@@
It is important to emphasize that this regularization technique is specifically designed to complement our curriculum training pipeline. For the regularization to be effective, the frozen teacher $T_{\mathrm{dn}}$ and the trainable teacher $T$ must have alignment in their output embeddings at the beginning of the restart training. This is accomplished by assigning them identical weights. Using a different frozen teacher, even if trained on the same denoised dataset, would not provide meaningful regularization due to misalignments, as shown in Appendix Figure˜9. Like NC, NCT is also denoiser-free at downstream fine-tuning and inference.
@@ 4.1 Practical and Challenging Experimental Settings
Noise Addition. We follow ImageNet-C [19] to introduce three types of noise: Gaussian, Shot (Poisson), and Speckle noise, as these are among the most commonly encountered noise types in natural imaging. For the ImageNet-100 experiments, we use Gaussian noise ($\sigma=50,100,255$; SNR: 8.73 dB, 3.98 dB, 0.31 dB), Shot noise ($\lambda=10,3,1$; SNR: 8.49 dB, 4.11 dB, 0.52 dB), and Speckle noise ($\sigma=102,178.5,255$; SNR: 9.15 dB, 5.58 dB, 3.98 dB). For the ImageNet-1k experiments, we tested Gaussian noise at $\sigma=100$ and $\sigma=255$, with corresponding SNRs of 4.36 dB and 0.72 dB respectively. These SNRs are common values reported in real noisy medical imaging [41]. The example visualizations for each noise level are provided in Section˜D.2. The noises are added to raw images before preprocessing (e.g., resizing) to better reflect practical scenarios such as device or environmental noise. Consequently, while the initial noise level is fixed, the preprocessing steps introduce variations in actual noise levels that challenge the model’s ability to adapt effectively. Notably, we generate the noise dataset once prior to training and do not apply noise on-the-fly during data loading, which closely simulate real-world scenarios where one starts with a fixed noisy dataset.
Implementation Details. In ImageNet-100 experiments, we use ViT-S/16 as the base architecture for DINOv2, and unless otherwise mentioned, train it for 200 epochs with a batch size of 40 using the AdamW optimizer. The base learning rate is $7.9\times 10^{-4}$ (scaled by the square root of batch size). The denoiser is trained for 100 epochs with a batch size of 4 using Adam. The linear probe uses a batch size of 128 and is trained for 12.5k steps. In ImageNet-1k, we adopt ViT-B/16 as the backbone and train DINOv2 for 100 epochs with a batch size of 512 and drop path rate of 0.1. The learning rate is $2.8\times 10^{-3}$. The denoiser is trained on a 100k subset with a batch size of 8 for 50 epochs (for Gaussian $\sigma=255$) and 90 epochs (for Gaussian $\sigma=100$). The linear probe uses batch size 1024 for 25k steps. Each training is done on a single RTX 4090 (~8h) or $\timesL4$0S GPUs (~65h). The strength parameters are determined via light parameter sweeps with the details provided in Section˜A.2.
@@ 4.2 Linear Probing Reveals Superior Noise Adaptation
We begin by evaluating how different methods adapt to various levels and types of noise using a 200-epoch training schedule on ImageNet-100. We train a linear classifier over a frozen DINOv2 backbone on the training set, and report classification performance over the validation set. We compare the classification performance of the DINOv2 backbone over four training strategies. (1) N2N + DINOv2: train the backbone on denoised images for 200 epochs and evaluate the model on denoised images (2) DINOv2 w/ NC: train the backbone on denoised images for 140 epochs, then restart training on noisy images for 60 epochs without regularization. Finally, evaluate the model on noisy images. The rationale for such partitions of epochs is provided in Appendix B.3. (3) DINOv2 w/ NCT: same as the DINOv2 w/ NC but with the regularization loss applied in the 60 epochs of noisy training. We only evaluate the NCT method on the four most noisy scenarios (lowest SNRs), i.e., Gaussian noise ($\sigma$=255, $\sigma$=100), Shot noise ($\lambda$=1), and Speckle noise ($\sigma$=255). (4) DINOv2: train the backbone on noisy images for 200 epochs, and evaluate the model on noisy images. Additionally, we train and evaluate the model on original clean images, referred to as Clean. Figure˜5 illustrates the classification accuracy curves.
As shown in Figure˜5, our proposed method (green) significantly outperforms the backbone trained solely on noisy images (red) across all noise types and noise levels. At moderate noise levels, i.e., the first two columns, the accuracy of our method closely matches and even exceeds the backbone that is trained and evaluated on explicitly denoised data (shown in orange lines). Only at extreme noise levels, i.e., the last column, the gap between DINOv2 w/ NC and N2N + DINOv2 widens, primarily due to the significant loss of the original image signal. For instance, as shown in Section˜D.2, with Gaussian noise at $\sigma=255$, many images become virtually unrecognizable to the naked eye. Hence, regularization is introduced to guide the model in these extremely noisy scenarios. As illustrated in Figure˜5, regularization DINOv2 w/ NCT improves the model’s performance at all four scenarios. The improvement amplifies as the gap between N2N + DINOv2 and DINOv2 w/ NC widens, with a substantial improvement of 6.7% over NC achieved at Gaussian noise with $\sigma=255$. However, when N2N + DINOv2 is on par with DINOv2 w/ NC, the regularization provides diminishing benefit, as demonstrated with Gaussian noise at $\sigma=100$. This is expected since the guidance becomes redundant when its representation quality is comparable to or lower than that of the original model.
@@
(a) Evolution of linear probing accuracies on ImageNet-100 under extended training durations. Relative ranking among methods remain consistent across epochs. See visualization in Figure˜2.
Method Clean
Gaussian-100
SNR: 4.36 dB
Gaussian-255
SNR: 0.72 dB
@@ 4.3 Scaling Improves Performance Across Datasets
Longer Training Duration. We further improve the performance of our methods by extending the training on ImageNet-100 to longer epochs. For 500-epoch configuration, we restart the training from the 200-epoch denoised checkpoint (i.e., N2N + DINOv2) as described in Section˜4.2; for the 1000-epoch configuration, we restart the training from the 500-epoch denoised checkpoint. As shown in Table˜1(a) and Figure˜2, accuracies increase as epoch increases and the relative ranking among methods remains consistent across training durations: N2N + DINOv2 achieves the highest accuracy, followed closely by DINOv2 w/ NC or NCT, while the baseline DINOv2 remains the lowest. This demonstrates the scalability of the proposed methods. Interestingly, we observe that DINOv2 w/ NCT consistently converges to or outperforms its anchor’s accuracy, as indicated by $\dagger$ and $\ddagger$ in Table˜1(a), demonstrating that the regularization objective successfully guides the model toward robust representations. Notably, the performance gap between DINOv2 and the other methods narrows as training progresses, suggesting that DINOv2 is sensitive to noise and benefits significantly from longer training. In contrast, our noise curriculum and regularization approach greatly accelerates convergence, achieving similar or better performance in roughly half the training time.
@@
Larger Dataset. In our controlled ImageNet-1k experiment, we constrain the DINOv2 backbone to undergo a total of 100 training epochs to enable fair comparison across methods. DINOv2 w/ NC: train the backbone on denoised images for 30 epochs followed by 70 epochs of training on noisy images. DINOv2 w/ NCT: While we find that restarting at 30 epochs yields the best performance for DINOv2 w/ NC, the denoised teacher at this early stage lacks sufficiently strong representations to serve as an effective regularizer. To address this, we continue training the denoised teacher to 100 epochs, then freeze it for regularizing the DINOv2 backbone for 70 epochs of noisy training. Since the frozen teacher and the DINOv2 backbone are derived from the same training run and pass through the same 30-epoch state, alignment is preserved. This allows us to meet the 100-epoch constraint while still leveraging a stronger teacher. Table˜1(b) shows that DINOv2 w/ NCT improves over the DINOv2 baseline with substantial absolute gains of $1.4\%$ and $4.8\%$ at noise levels $\sigma=100$ and $\sigma=255$, respectively. DINOv2 w/ NC also achieves a strong improvement of $2.5\%$ under the more challenging $\sigma=255$ setting. The relatively smaller gain of DINOv2 w/ NC at $\sigma=100$ is due to the fact that, in large-scale settings with moderate noise, longer training alone significantly improves DINOv2’s robustness. Since DINOv2 w/ NC allocates a portion of the training budget to denoised pretraining, it shortens the noisy training phase, slightly limiting its benefit. Overall, while scaling training on large datasets improves DINOv2’s noise adaptation under moderate noise, denoised pretraining and regularization remain essential under high noise to maintain strong performance. As an additional baseline, we also evaluate DINOv2 trained exclusively on clean images and directly tested on noisy inputs. The results presented in Section˜C.1 confirm that clean-pretrained model suffers severe accuracy drop, which underscores the importance of our noise-aware training.